1.1.11 Express the confidence interval (0.054,0.152) in the form of p-E<p<P+E. [<p< (Type integers or decimals.)
Answer #1

The midpoint of the interval is p

\widehat{p}=\frac{0.054+0.152}{2}=0.103

Next, the error term E is the distance from p to the end of the interval.

E=0.152-0.103=0.049

Finally, write the final form:

\widehat{p}\pm E=0.103\pm 0.049

\widehat{p}- E<p<\widehat{p}+E:0.103- 0.049<p<0.103+0.049

Hope it helps! Please give a thumbs up if it does and feel free to ask any doubts in the comments.